### v1

# New entropy, thermodynamics of apparent horizon and cosmology

Abstract
Here, we consider new nonadditive entropy of the apparent horizon $S_{K}=S_{BH}/(1+\gamma S_{BH})$ with $S_{BH}$ being the Bekenstein–Hawking entropy. This is an alternative of the Rényi and Tsallis entropies, that allow us by utilising the holographic principle to develop of a new model of holographic dark energy. When $\gamma\rightarrow 0$ our entropy becomes the Bekenstein–Hawking entropy $S_{BH}$. The generalised Friedmann’s equations for Friedmann–Lemaître–Robertson–Walker (FLRW) space-time for the barotropic matter fluid with $p=w\rho$ were obtained. We compute the dark energy pressure $p_{D}$, density energy $\rho_{D}$ and the deceleration parameter corresponding to our model. From the second modified Friedmann’s equation a dynamical cosmological constant was obtained. We show that at some model parameters $w$ and $\beta$ there are two phases, universe acceleration and deceleration or the phase of the eternal inflation. Thus, our model, by virtue of the holographic principle, can describe the universe inflation and late time of the universe acceleration. The holographic dark energy model with the generalised entropy of the apparent horizon can be of interest for new cosmology.
I Introduction
To describe the current accelerating universe one can introduce in the Einstein equations the cosmological constant. It plays the role of dark energy leading to standard cosmology. Another way to describe acceleration of universe is to explore thermodynamics of apparent horizon Akbar ; Cai ; Cai0 ; Paranjape ; Sheykhi ; Cai1 ; Wang ; Jamil ; Gim ; Fan ; Agostino ; Sanchez because there is a correspondence between gravity and thermodynamics. This is based on the fact that as was shown in Refs. Bekenstein ; Hawking that in black holes, entropy is connected with the horizon area and temperature is linked with the surface gravity Jacobson ; Padmanabhan ; Padmanabhan1 ; Hayward . Also, Friedmann’s equations may be obtained with the help of the first law of apparent horizon thermodynamics. As a result, the apparent horizon of the FLRW space-time is represented as a thermodynamic system and it is a causal boundary Hayward ; Hayward1 ; Bak , and on this boundary the thermodynamics laws are satisfied Akbar ; Cai0 . In flat FLRW universe the apparent horizon is equal to the Hubble horizon. Different entropies Tsallis ; Ren ; Renyi ; Jahromi ; Barrow ; Kaniadakis ; Mejrhit ; Majhi were introduced, due to the long-range nature of gravity, that lead to the generalized Friedmann equations. Motivated by success of this approach, it is of interest to consider other entropies to describe accelerating universe. Because entropies are sources of holographic energy densities they can describe the dark energy of the universe Pavon ; Landim . In this letter, we consider new apparent horizon entropy $S_{K}=S_{BH}/(1+\gamma S_{BH})$ where $S_{BH}$ is the Bekenstein–Hawking entropy and it is a non-extensive entropy measure. When the Bekenstein–Hawking entropy becomes zero our entropy $S_{K}$ also vanishes, $S_{K}$ is the monotonically increasing function of $S_{BH}$ and is positive. As $\gamma\rightarrow 0$ we come to the Bekenstein–Hawking entropy. Here, we consider the case with equation of state (EoS) for barotropic perfect fluid, $p=w\rho$. We will obtain modified Friedmann’s equations with dynamical cosmological constant that leads to dark energy and the universe inflation. Our new holographic model with the non-additive entropy and non-extensive feature of space-time shows the possibility of the universe to accelerate in accordance with observations.
We assume units with $\hbar=c=k_{B}=1$.
II New entropy
We start with new entropy
$S_{K}=-\sum_{i=1}^{W}\frac{p_{i}\ln{p_{i}}}{1-\gamma\ln{p_{i}}}),$ (1)
where $W$ is a number of states. Each state possesses a probability $p_{i}$ with the probability distribution $\{p_{i}\}$ and $\gamma$ is a free parameter. In Eq. (2.1) the summation is performed over all possible system microstates. When $\gamma=0$ entropy (2.1) is converted into the Gibbs entropy
$S_{G}=-\sum_{i=1}^{W}p_{i}\ln(p_{i}).$ (2)
If one assumes that each microstate is populated with equal probability, then $1/p_{i}=W$ ($i=1,2,...,W$) and Eq. (2.2) is converted into the Boltzmann entropy $S_{B}=\ln(W)$. Making use of $1/p_{i}=W$, we obtain from Eq. (2.1)
$S_{K}=\frac{\ln(W)}{1+\gamma\ln(W)}.$ (3)
The Bekenstein–Hawking entropy $S_{BH}=\ln(W)$. Then from Eq. (2.3) one finds
$S_{K}=\frac{S_{BH}}{1+\gamma S_{BH}}.$ (4)
Making use of Eq. (2.4), at $\gamma=0$, we obtain the Bekenstein–Hawking entropy $S_{BH}$. If A and B are two probabilistically independent systems, we have $p^{A+B}_{ij}=p^{A}_{i}p^{B}_{j}$ and one finds nonadditive entropy, $S_{K}(A+B)\neq S_{K}(A)+S_{K}(B)$.
III Apparent horizon thermodynamics
Let us consider the FLRW flat universe which is described by the metric
$ds^{2}=-dt^{2}+a(t)^{2}(dr^{2}+r^{2}d\Omega_{2}^{2}).$ (5)
Here, $a(t)$ is a scale factor and $d\Omega_{2}^{2}$ is the line element of an 2-dimensional unit sphere. The radius of the apparent horizon $R_{h}=a(t)r$, in the FLRW universe, is given by
$R_{h}=\frac{1}{H},$ (6)
where the Hubble parameter of the universe is $H=\dot{a}(t)/a(t)$, with dot over $a(t)$ being the derivative with respect to the cosmological time $t$. The first law of apparent horizon thermodynamics is formulated as
$dE=-T_{h}dS_{h}+WdV_{h},$ (7)
where $W$ is the work density and $E$ is the total energy inside the space which is given by
$E=\rho V_{h}=\frac{4\pi}{3}\rho R_{h}^{3}.$ (8)
The change of the energy inside the apparent horizon is $-dE$. In Eq. (3.4) $\rho$ means the energy density of matter fields and the work density is Hayward ; Hayward1 ; Bak
$W=-\frac{1}{2}\mbox{Tr}(T^{\mu\nu})=\frac{1}{2}(\rho-p),$ (9)
and $p$ is the matter pressure. In our case the horizon entropy is $S_{h}=S_{K}$. The apparent horizon temperature is defined as
$T_{h}=\frac{H}{2\pi}\left|1+\frac{\dot{H}}{2H^{2}}\right|.$ (10)
Making use of first law of apparent horizon thermodynamics (3.3), and by combining Eqs. (3.4) and (3.5) with (3.6) , one finds
$\frac{H}{2\pi}\left|1+\frac{\dot{H}}{2H^{2}}\right|dS_{h}=-\frac{4\pi}{3H^{3}}d\rho+\frac{2\pi(\rho+p)}{H^{4}}dH.$ (11)
With the help of the continuity equation
$\dot{\rho}=-3H(\rho+p),$ (12)
and Eq. (3.7) we obtain
$\frac{H}{2\pi}\left|1+\frac{\dot{H}}{2H^{2}}\right|\dot{S}_{h}=-\frac{4\pi\dot{\rho}}{3H^{3}}\left(1+\frac{\dot{H}}{2H^{2}}\right).$ (13)
IV Generalized Friedmann’s equations
Utilizing Eqs. (3.8), (3.9) and assuming that $01+\dot{H}/(2H^{2})>$ 0 we obtain
$\frac{H}{2\pi}\dot{S}_{h}=\frac{4\pi(\rho+p)}{H^{2}}.$ (14)
By using our entropy function
$S_{h}=S_{K}=\frac{S_{BH}}{1+\gamma S_{BH}},$ (15)
where $S_{BH}=\pi R^{2}_{h}/G=\pi/(GH^{2})$, we obtain from Eq. (4.1) the generalized Friedmann equation
$\frac{\dot{H}}{(1+\gamma\pi/(GH^{2}))^{2}}=-4\pi G(\rho+p).$ (16)
At $\gamma=0$ in Eq. (4.3), we come to the first Friedmann equation for flat universe within Einstein gravity. Making use of Eq. (3.8) and after integration of Eq. (4.3), we find the second generalized Friedmann equation
$H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)=\frac{8\pi G}{3}\rho,$ (17)
where $b=\pi\gamma/G$ and we have used the integration constant $C=2b\ln(b)$. At $\gamma=0$ ($b=0$) one obtains from Eq. (4.4) the Friedmann equation of general relativity. We can represent Eq. (4.4) as follows:
$H^{2}=\frac{8\pi G}{3}\rho+\frac{\Lambda_{eff}}{3}.$ (18)
with
$\Lambda_{eff}=\frac{3b^{2}}{H^{2}+b}+6b\ln\left(\frac{H^{2}+b}{b}\right).$ (19)
The $\Lambda_{eff}$ represents the dynamical (effective) cosmological constant. We depicted the dynamical cosmological constant $\Lambda_{eff}$ versus $H$ at $b=1,2,3$ in Fig. 1.
Figure 1: The function $\Lambda_{eff}$ versus $H$ at $b=\pi\gamma/G=1,2,3$. Figure 1 shows that $\Lambda_{eff}$ increases as $b$ increases. At $H\rightarrow 0$ ($R_{h}\rightarrow\infty$) we have $\Lambda_{eff}\rightarrow 3b$.
According to Fig. 1 when $b$ increases, at fixed $H$, the dynamical cosmological constant $\Lambda_{eff}$ also increases and $\lim_{H\rightarrow 0}\Lambda_{eff}=3b$. By virtue of Eq. (4.5) one finds the density of dark energy
$\rho_{D}=\frac{3b}{8\pi G}\left[\frac{b}{H^{2}+b}+2\ln\left(\frac{H^{2}+b}{b}\right)\right].$ (20)
Defining the normalized density parameters $\Omega_{m}=\rho/(3M_{P}^{2}H^{2})$ and $\Omega_{D}=\rho_{D}/(3M_{P}^{2}H^{2})$, where $M_{P}=1/\sqrt{8\pi G}$ is the reduced Plank mass, one obtains from Eqs. (4.4) and (4.7) that $\Omega_{m}+\Omega_{D}=1$. From Eq. (4.4) one obtains the normalized density for the matter $w=0$
$\Omega_{m}=1-\frac{b^{2}}{(H^{2}H^{2}+b)}-2\frac{b}{H^{2}}\ln\left(\frac{H^{2}+b}{b}\right).$
The plot of $\Omega_{m}$ versus $H$ is depicted in Fig. 2.
Figure 2: The function $\Omega_{m}$ versus $H$ at $b==0.15,0.2,0.3$. Figure 2 shows that $\Omega_{m}$ increases as $b$ increases. As $H\rightarrow\infty$ ($R_{h}\rightarrow 0$) we have $\Omega_{m}\rightarrow 1$.
According to Fig. 2 for the current era $\Omega_{m}=0.26$ that can be realized for some parameter $b$. We will imply here that there is no mutual interaction between various components of the cosmos, from ordinary conservation law, one obtains
$p_{D}=-\frac{\dot{\rho}_{D}}{3H}-\rho_{D}.$ (21)
By virtue of Eqs. (4.7) and (4.8) we find the pressure corresponding to the dark energy
$p_{D}=-\frac{b(b+2H^{2})\dot{H}}{4\pi G(b+H^{2})^{2}}-$
$\frac{3b}{8\pi G}\left[\frac{b}{H^{2}+b}+2\ln\left(\frac{H^{2}+b}{b}\right)\right].$ (22)
Making use of Eqs, (4.3), (4.4) and (4.9) we obtain
$p_{D}=\frac{3b(b+2H^{2})(1+w)}{8\pi GH^{4}}\biggl{[}H^{2}-\frac{b}{H^{2}+b}-$
$2\ln\left(\frac{H^{2}+b}{b}\right)\biggr{]}-\frac{3b}{8\pi G}\left[\frac{b}{H^{2}+b}+2\ln\left(\frac{H^{2}+b}{b}\right)\right].$ (23)
With the help of Eqs. (4.7) and (4.10) one finds EoS for dark energy
$w_{D}=\frac{p_{D}}{\rho_{D}}=$
$\frac{(b+2H^{2})(1+w)}{H^{4}}\left[\frac{H^{2}(H^{2}+b)}{b+2(H^{2}+b)\ln\left(\frac{H^{2}+b}{b}\right)}-1\right]-1.$ (24)
The plot of $w_{D}$ versus $H$ is depicted in Fig. 3.
Figure 3: Left panel: The function $w_{D}$ versus $H$ at $b=1$, $w=0,1/3,2/3$, $G=1$, $\rho_{0}=a_{0}=1$. According to Fig. 3 the $w_{D}$ increases as EoS parameter for the matter $w$ increases, at fixed $H$. At large $H$ EoS parameter for dark energy $w_{D}$ approaches to $-1$. Right panel: In accordance with figure when parameter $b$ increases, at fixed $H$, EoS parameter for dark energy $w_{D}$ also increases and $\lim_{H\rightarrow\infty}w_{D}=-1$.
In accordance with Fig. 3 (Left panel) at $b=1$ and $w=0,1/3,2/3$ when $w$ increases, at fixed $H$, EoS parameter for dark energy $w_{D}$ also increases. Figure 3 shows (Right panel) that at $w=1$ and $b=2,3,4$ if $b$ increases, at fixed $H$, $w_{D}$ aolso increases. It follows from Eq. (4.11) that $\lim_{H\rightarrow\infty}w_{D}=-1$. Thus, dynamical cosmological constant leads to dark energy EoS $w_{D}=-1$ at large Habble parameter $H$.
The second law of thermodynamics requires that $\dot{S}_{K}\geq 0$. Then from Eq. (3.2) one obtains $\dot{S}_{BH}/(1+\gamma S_{BH})^{2}\geq 0$ and this leads to $\dot{S}_{BH}=-2\pi\dot{H}/(GH^{3})\geq 0$. As a result, we have the same requirement as for the Bekenstein–Hawking entropy. For positive Hubble parameter one obtains $\dot{H}\leq 0$. Then from Eq. (3.3) we find $\rho+p\geq 0$ and for $\rho>0$ the EoS parameter $w\geq-1$. It is convenient to use the redshift $z=a_{0}/a(t)-1$, where $a_{0}$ corresponds to a scale factor at the current time. With the aid of continuity equation (3.8) and EoS $p=w\rho$ one obtains the matter density energy
$\rho=\rho_{0}\left(\frac{1+z}{a_{0}}\right)^{3(1+w)},$ (25)
with $\rho_{0}$ being the density energy of matter at the present time. From Eqs. (4.4) and (4.12) we find equation as follows:
$H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)=\frac{8\pi G\rho_{0}}{3}\left(\frac{1+z}{a_{0}}\right)^{3(1+w)}.$ (26)
The Hubble parameter versus redshift $z$ for $G=1$, $\rho_{0}=a_{0}=1$ is plotted in Fig. 4.
Figure 4: Left panel: The function $H$ versus $z$ at $b=0.1$, $w=1/3,0,-0.1$, $G=1$, $\rho_{0}=a_{0}=1$. According to Fig. 4 $H$ increases as $z$ increases. At fixed $H$, when EoS parameter for the matter $w$ increases the redshift $z$ decreases. Right panel: In accordance with figure if parameter $b$ increases at fixed $z$ the Habble parameter also increases.
When redshift $z$ increases the Habble parameter also increases. According to Fig. 4 (Left panel) if parameter $w$ increases the redshift $z$ decreases. Figure 4 (Right panel) shows that when parameter $b$ increases at fixed z the Habble parameter also increases The deceleration parameter is given by
$q=-\frac{\ddot{a}a}{\dot{a}^{2}}=-1-\frac{\dot{H}}{H^{2}}.$ (27)
If $q<0$ we have the acceleration phase of the universe and when $q>0$ the universe decelerates. Making use of Eqs. (4.3), (4.12) and (4.14) we find
$q=\frac{4\pi G\rho_{0}(1+w)(H^{2}+b)^{2}}{H^{6}}\left(\frac{1+z}{a_{0}}\right)^{3(1+w)}-1.$ (28)
Equation (4.15) defines the dependence of the deceleration parameter $q$ on redshift $z$. By virtue of Eqs. (4.13) and (4.15) we obtain the function of deceleration parameter $q$ on $H$
$q=-1+$
$\frac{3(1+w)(H^{2}+b)^{2}}{2H^{6}}\left(H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)\right).$ (29)
We depicted the deceleration parameter $q$ versus the Hubble parameter $H$ for $\rho_{0}=a_{0}=1$ in Fig. 5.
Figure 5: According to Fig. 5 when deceleration parameter $q$ increases the Habble parameter $H$ also increases. Left panel: The deceleration parameter $q$ versus Habble parameter $H$ at $b=1$, $w=0,1/3$, 2/3, $\rho_{0}=a_{0}=1$. When EoS parameter for the matter $w$ increases, at fixed $b$ and $H$, the deceleration parameter $q$ also increases. There are two phases: universe acceleration $q<0$ and deceleration $q>0$. Right panel: in accordance with figure, when parameter $b$ increases, at fixed $w$ and $H$, the deceleration parameter $q$ decreases. Here we have two phases, universe acceleration and deceleration.
Figure 5 shows that there are two phases: universe acceleration and deceleration. Making use of Eq. (4.16) we find the asymptotic
$\lim_{H\rightarrow\infty}q=\frac{3w+1}{2}.$ (30)
Equation (4.17) shows that the asymptotic of the deceleration parameter as $H\rightarrow\infty$ ($R_{h}\rightarrow 0$) does not depend on the entropy parameter $\gamma$. It follows from Eq. (4.17) that when $w>-1/3$ ($q>0$) we have two phases: universe acceleration and deceleration. If $w<-1/3$ the eternal inflation takes place. By virtue of Eq. (4.13) we obtain the redshift $z$ as a function of Habble parameter
$z=-1+$
$a_{0}\left[\frac{3}{8\pi\rho_{0}G}\left(H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)\right)\right]^{1/(3(1+w))}.$ (31)
From Eqs. (4.16) and (4.18) we found the approximate real and positive solutions for the transition redshifts $z_{t}$ (from acceleration to deceleration phases) at $q=0$, $G=1$, $w=0$ (for the matter dominated era) presented in Table 1.
Table 1: The approximate solutions to Eqs. (3.16) and (3.18) for the transition redshifts $q=0$ at $G=1$, $a_{0}=\rho_{0}=1$, $w=0$.
$b$ 1.5 2 2.5 3 3.5 4 4.5 5 5.5 6 6.5
$H$ 4.26 4.92 5.50 6.03 6.51 6.96 7.38 7.78 8.16 8.53 8.88
$z_{t}$ 0.07 0.18 0.27 0.35 0.42 0.49 0.55 0.60 0.65 0.70 0.75
According to Table I when parameter $b$ increases (at fixed $w$) the Hubble parameter $H$ and reshift $z$ also increase for a point $q=0$ between universe acceleration and deceleration. From Eqs. (4.16) and (4.18) for the current era $z=0$, we find solutions for the Habble parameter $H$ and deceleration parameter $q$ for varies $b$, given in Table II.
Table 2: The solutions (approximate) to Eqs. (3.16) and (3.18) for the current era $z=0$ at $G=1$, $a_{0}=\rho_{0}=1$, $w=-2/3$.
$b$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
$H$ 3.05 3.15 3.25 3.33 3.40 3.47 3.54 3.60 3.66
$q$0.54 -0.56 -0.58 -0.59 -0.61 -0.62 -0.63 -0.64 -0.65
In Table II the deceleration parameter $q$ is negative, and therefore, we have here the acceleration phase in the current time. The deceleration parameter at the current era is $q_{0}\approx-0.6$ Roos corresponding to parameter $b\approx 0.5$. By virtue of (3.16) we plotted the Habble parameter $H$ versus redshift $z$ in Fig. 6.
Figure 6: Left panel: The Habble parameter $H$ versus $z$ at $b=0.1$, $w=-0.3,1/3$, 2/3, $G=1$, $\rho_{0}=a_{0}=1$. Figure 6 shows that when $z$ increases, $H$ also increases. When EoS parameter $w$ increases, at fixed $b$, the Habble parameter $H$ decreases. Right panel: When parameter $b$ increases, at fixed $w$, the Habble parameter $H$ also increases.
In accordance with Fig. 6, when redshift $z$ increases, the Habble parameter $H$ also increases. When EoS parameter $w$ increases, at fixed $b$, the Habble parameter $H$ also increases. Figure 6 shows that if parameter $b$ increases, at fixed $w$, the Habble parameter $H$ also increases. Making use of Eqs. (4.16) and (4.18) we depicted the deceleration parameter $q$ versus redshift $z$ in Fig. 7.
Figure 7: Left panel: The deceleration parameter $q$ versus $z$ at $b=0.1$, $w=-0.3,0$, 1/3, $G=1$, $\rho_{0}=a_{0}=1$. In accordance with Figure 7, when $z$ increases the deceleration parameter $q$ decreases. When EoS parameter for the matter $w$ increases, at fixed $b$, the deceleration parameter $q$ also increases. At $w=-0.3$ there is two phases, universe acceleration $q<0$ and deceleration $q>0$ and at $w=0$ and $1/3$ we have only the deceleration phase. Right panel: According to Fig. 5 when parameter $b$ increases, at fixed $w$, the deceleration parameter $q$ also increases. We have here two phases: acceleration and deceleration for $b=0.4$ and $0.6$. For $b=0.2$ one has only the deceleration phase.
Figure 7 shows that when $z$ increases $q$ also increases. When EoS parameter for the matter $w$ increases, at fixed $b$ (Left panel), the deceleration parameter $q$ also increases. At $w=-0.3$ there are two phases, universe acceleration $q<0$ and deceleration $q>0$ but at $w=0$ and $w=1/3$ one has only the deceleration phase. According to Fig. 5 (Right panel), when parameter $b$ increases at fixed $w$ the deceleration parameter $q$ decreases. One has here two phases, acceleration and deceleration for $b=0.4$ and $b=0.6$ but for $b=0.2$ we have the deceleration phase.
V Conclusion
Considering a novel entropy $S_{K}=S_{BH}/(1+\gamma S_{BH})$ we have a property similar to the Bekenstein–Hawking entropy $S_{BH}$. By using our entropy and by employing the holographic principle, we have obtained a dark energy model. The $S_{K}$ vanishes when the apparent horizon radius $R_{h}$ is zero and it increases monotonically when the apparent horizon radius $R_{h}$ increases. The barotropic perfect fluid and flat FLRW universe were studied. We have obtained, from first law of apparent horizon thermodynamics, the modified Friedmann’s equations. The second Friedmann’s equation possesses an addition term corresponding to the density of dark energy which can be treated as a dynamical cosmological constant. Implying that there is no interaction between various components of cosmos, and the dark energy density and pressure $\rho_{D}$ and $p_{D}$ obeys ordinary conservation law, we computed $\rho_{D}$ and $p_{D}$. The EoS parameter for dark energy $w_{D}=p_{D}/\rho_{D}$ was calculated and it was shown that $\lim_{H\rightarrow\infty}w_{D}=-1$. The universe can have two phases, acceleration and deceleration, that is due to holographic dark energy. It should be noted that Barrow and Tsallis entropies also lead to cosmology which ia due to Einsten’s equations with the dynamical cosmological constant Gennaro . We computed the deceleration parameter that shows the possibility, for some model parameters, to describe inflation and acceleration at current era. The transition redshifts (when $q=0$) was calculated (given in Table I) for some parameters. We show (see Table II) that at $b\approx 0.5$ and $w=-2/3$ we have the deceleration parameter $q_{0}\approx-0.6$ which was observed in universe at the current era Roos . Thus, our approach, based on new entropy and leading to modified Friedmann equations, can description universe inflation and late time universe acceleration. It is worth mentioning that the results obtained here are similar to those in another model with different entropy Kruglov .
Data availability statement: No data was used for the research described in the article.